# HATS-38 b and WASP-139 b join a growing group of eccentric hot Neptunes on polar orbits [^1] Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programs ID 111.24VT.001, 111.24VT.002, and 112.25W1.001

### v1

@@ Abstract
We constrain the sky-projected obliquities of two low-density hot Neptune planets, HATS-38 b and WASP-139 b, orbiting nearby G and K stars using Rossiter-McLaughlin (RM) observations with VLT/ESPRESSO, yielding $\lambda=-108_{-16}^{+11}$ deg and $-85.6_{-4.2}^{+7.7}$ deg, respectively. To model the RM effect, we use a new publicly available code, ironman, which is capable of jointly fitting transit photometry, Keplerian radial velocities, and RM effects. The two planets have residual eccentricities ($e=0.112_{-0.070}^{+0.072}$, and $0.103_{-0.041}^{+0.050}$, respectively), and together with the obliquity constraints, we show that they join a growing group of eccentric hot and low-density Neptunes on polar orbits. We use long-term radial velocities to rule out companions with masses $\sim 0.3-50$ $M_{J}$ within $\sim$10 au. We show that the orbital architectures of the two Neptunes disfavor an origin from primordial disk misalignment and/or in-situ secular interactions with distant companions and instead favor high-eccentricity migration from $\gtrsim 2$ au driven by a distant companion. Finally, we performed a hierarchical Bayesian modeling of the true obliquity distribution of Neptunes and found suggestive evidence for a higher preponderance of polar orbits of hot Neptunes compared to Jupiters. However, we note that the exact distribution is sensitive to the choice of priors, highlighting the need for additional obliquity measurements of Neptunes to robustly compare the hot Neptune obliquity distribution to Jupiters.
@@ 1 Introduction
However, the signature of the dynamical evolution is the true 3D obliquity $\psi$, not its sky projection $\lambda$. The true obliquity $\psi$, can be estimated from $\lambda$ measured by the RM effect through combining it with constraints on the orbital inclination, $i$ constrained by the shape of a planetary transit, and the stellar inclination, $i_{\star}$, which can be constrained by measuring the stellar rotation period ($P_{\rm rot}$), the projected rotational velocity ($v\sin i_{\star}$), and the radius of the host star (e.g., Hirano et al., 2014; Morton & Winn, 2014; Masuda & Winn, 2020).
@@
Thanks to recent observational efforts, the current sample of planets with measured obliquity shows a growing number of smaller planets including hot Neptunes ($5010<M_{p}/M_{\oplus}<$ 50 or $62<R_{p}/R_{\oplus}<$ 6) in the ‘hot Neptune desert’, a region of the period-radius diagram where Neptune-sized planets are notably rare (Mazeh et al., 2016) likely due to intense stellar radiation causing atmospheric erosion. Interestingly, many hot Neptunes are observed to be in close-to-polar orbits and have evaporating atmospheres, including HAT-P-11 b (Sanchis-Ojeda & Winn, 2011), GJ 436 b (Bourrier et al., 2018, 2022), WASP-107 b (Dai & Winn, 2017; Rubenzahl et al., 2021), and GJ 3470 b (Stefansson et al., 2022), suggesting a possible link between the Neptune desert and the dynamical history of low-mass gas giants. Theoretical scenarios that are capable of explaining the polar orbits of these planets require the presence of massive distant companions (e.g., Petrovich et al., 2020). Recently, outer companions have been confirmed in the HAT-P-11 (Yee et al., 2018) and WASP-107 (Piaulet et al., 2021) systems, suggesting that the transiting planets arrived at their current polar orbits through dynamical interactions (see Lu et al., 2024; Yu & Dai, 2024). As for the atmospheres, the orbital eccentricities may be large enough so that tidal heating can inflate these planets (e.g., by a factor of $\simeq 1.5-1.7$ for GJ 3470b; Stefansson et al., 2022), possibly accounting for their evaporation. Nevertheless, the sample of hot Neptunes with measured obliquities is still small, and more measurements are necessary to test these emerging trends and asses the underlying mechanism behind the formation of this transiting population.
In this work, we present ESPRESSO (Pepe et al., 2021) observations of the RM effect of the two hot Neptunes HATS-38 b (Jordán et al., 2020) and WASP-139 b (Hellier et al., 2017). HATS-38 b is a hot Neptune in the middle of the hot Neptune desert orbiting a G-dwarf host star, while WASP-139 b is a super-Neptune orbiting a K-star. The precise in-transit spectroscopic observations with ESPRESSO reveal that both planets are in nearly-polar and eccentric orbits, continuing with the observed trends for low-density gas giants. We also performed a long-term radial velocity (RV) monitoring of both stars in order to constrain the presence of distant planetary companions in the systems.
@@
CHAT 2017 Feb 05 $r$ filter Jordán et al. (2020)
@@ 2.3 Observatoire Moana Photometry
Simultaneously with ESPRESSO observations, we observed the transit of HATS-38 b using the station of the Observatoire Moana located in El Sauce (ES) Observatory in Chile. Observatoire Moana is a global network of small-aperture robotic optical telescopes. The ES station consists of a 0.6 m CDK robotic telescope coupled to an Andor iKon-L 936 deep depletion 2k $\times$ 2k CCD with a scale of 0.67′′ per pixel. For these observations, we used a Sloan $r^{\prime}$ filter, and the exposure time was set to 80 s. The MOANA/ES light curve is shown in panel h of Figure 1, along with the best model.
@@ 3 Stellar Parameters
The second step consists of obtaining the physical parameters of the star by using publicly available broad-band magnitudes of the star and comparing them with those produced by different PARSEC stellar evolutionary models (Bressan et al., 2012) by taking into account the distance to the star computed from the Gaia DR2 (Gaia Collaboration et al., 2018) parallax. This procedure delivers a new value of $\log{g}$ that is held fixed in a new run of zaspe. We iterate between the two procedures until reaching convergence, which happens when two consecutive zaspe runs deliver the same values of $T_{\rm eff}$ and [Fe/H]222these parameters come from a grid search, so that for small changes in $\log{g}$ ($\Delta\log{g}<0.01$) the model with the best matching $T_{\rm eff}$ and [Fe/H] will be the same. . The obtained parameters for each star are presented in Table 2.
@@ 5 Joint fit
Figure 3: Sky-projected obliquities of Neptunes ($5010<M_{p}/M_{\oplus}<$ 50 or $62<R_{p}/R_{\oplus}<$ 6) as a function of stellar effective temperature (blue). Neptunes on compact multiplanet systems are shown as squares, while isolated Neptunes are shown as circles. Jupiters (not Neptunes with $R_{p}/R_{\oplus}>6$) and super-Earths (not Neptunes with $R_{p}/R_{\oplus}<2$) from TEPCat (Southworth, 2011) are shown in orange and grey, respectively.
With ironman, we analyzed both systems, including all the observations presented in Section 2. For this analysis, we only considered the detrended TESS light curves (see Section 4) close to the transits, in windows of $\sim 6$ hours around the transit midpoint, to reduce the computational cost. To account for the instrumentation offsets and systematics, we included independent RV offsets and log-uniform jitter terms for each instrument. We placed uninformative priors for almost all parameters, except for the orbital periods and times of mid-transit, which had been constrained in Section 4, and for the stellar densities that were constrained in Section 3. Likewise for $\beta$ where we considered an instrumental broadening of 2.15 km/s because of the ESPRESSO resolution, and we used the macroturbulence law from Valenti & Fischer (2005), which yields a broadening of $\sim 4.15$ and $\sim 4.8$ km/s for $T_{\rm eff}=5662$ and $5233$ K, respectively. We added the instrumental and macroturbulence broadening in quadrature to set our priors for each target respectively, with an uncertainty of 2 km/s. The resulting posteriors of our fits for HATS-38 b and WASP-139 b are shown in Table 3.
Figures 1 and 2 show the different datasets together with the best-fit model. We found that both planets lie in a nearly-polar orbit, with sky-projected obliquities $\lambda=-108_{-16}^{+11}$ and $-85.6_{-4.2}^{+7.7}$ deg for HATS-38 b and WASP-139 b, respectively. Also, we found that both planets are consistent with having eccentric orbits. This was not noted in Jordán et al. (2020) and Hellier et al. (2017), where they imposed circular orbits. We tested fitting circular orbits as well, but the eccentric fit had a higher evidence value, with a $\Delta\log{Z}>2$ in both cases, so it is the one we adopted. We also tried fitting a long-term linear RV slope to the data, but it resulted in being consistent with zero so we fixed it to that value. Overall, the rest of the parameters are consistent with the ones reported in Jordán et al. (2020) and Hellier et al. (2017).
@@ 6 Discussion
In this work, we have characterized the orbits HATS-38 b and WASP-139 b, and found that they have nearly polar sky-projected obliquities and residual non-zero eccentricities, thus joining the growing sample of isolated (without nearby companions within 4 times the orbital period) low-density Neptunes in polar and eccentric orbits that include HAT-P-11 b (Sanchis-Ojeda & Winn, 2011), WASP-107 b (Dai & Winn, 2017; Rubenzahl et al., 2021), GJ 436 b (Bourrier et al., 2018, 2022), and GJ 3470 b (Stefansson et al., 2022).
@@
$\beta_{\rm ESPRESSO}$ Intrinsic ESPRESSO stellar line width (km/s) $4.4_{-2.0}^{+1.9}$ $5.7_{-1.8}^{+1.8}$
@@ 6.1 A possible preponderance of polar Neptunes?
To better quantify the possible preponderance of polar Neptunes in the sample, using the two new measurements we have provided here, we used a hierarchical Bayesian model (e.g., Hogg et al., 2010) for the underlying obliquity distribution (e.g., Morton & Winn, 2014; Muñoz & Perets, 2018). In particular, we used the framework [^5] https://github.com/jiayindong/obliquity presented by Dong & Foreman-Mackey (2023), which models the underlying distribution of $\cos{\psi}$ across an exoplanet population using a mixture model of two Beta distributions (e.g., Gelman et al., 2014). This hierarchical Bayesian modeling framework allows the user to derive the true 3D obliquity distribution from observed sky-projected obliquities and has the capacity to capture anything from an isotropic distribution to a strongly bimodal population.
@@
Figure 4: Inferred stellar obliquity distribution of Neptunes (blue) and Jupiters (orange) with sky-projected stellar obliquity measurements. This inference was done following the methodology from Dong & Foreman-Mackey (2023) using the observed sky-projected obliquities available in the literature—combined with the two sky-projected obliquity measurements presented here. The lines and shaded regions represent the median and $1\sigma$ uncertainties of the inferred distributions. We show the results for a) a $\mathcal{N}(0,3)$ prior on the $\log{\kappa}$ parameter and b) a $\mathcal{U}(-4,10)$ prior on $\log{\kappa}$.
To derive the underlying $\cos{\psi}$ distribution, we ran two different hierarchical models with different priors on the compactness parameter $\kappa$: first, a model using informative priors on $\kappa$ following Dong & Foreman-Mackey (2023), and second, a fit using uninformative priors on $\kappa$. For the first model, we followed Dong & Foreman-Mackey (2023), and assumed an informative normal prior on $\log{\kappa}$ with a mean of 0 and a standard deviation of 3. Second, as an additional test—given the small size of the sample—we ran another model with an uninformative uniform prior between -4 and 10 for $\log{\kappa}$. In both cases, we used uniform priors on the location parameter $\mu$. The results of both of these model fits are shown in Table 4, and in Figure 4.
@@
The observed difference in the obliquity distribution for the sample of Neptunes and Jupiters—with the former showing a possible preference for a polar + aligned obliquity dichotomy, while the latter shows a preference for an aligned + isotropic dichotomy—likely suggests that hot Jupiters and hot Neptunes could have formed through different formation channels. However, the sample of Neptunes is small compared to the one of Jupiters (27 versus 171 systems), and the exact parameters of the bimodal distribution are different with different priors. According to Dong & Foreman-Mackey (2023), in their tests, prior distribution choices did not strongly affect the derived distributions for samples of more than $\sim 50$ systems. Therefore, more measurements of the stellar obliquity of Neptunes will be required to further confirm or rule out these results. Future efforts like this one, and programs like The Desert-Rim Exoplanets Atmosphere and Migration (DREAM, Bourrier et al., 2023), will be crucial for increasing the size of the sample and understanding the formation and evolution of the hot Neptunes.

### v2

# HATS-38 b and WASP-139 b join a growing group of hot Neptunes on polar orbits [^1] Based on observations made with ESO Telescopes at the La Silla Paranal Observatory under programs ID 111.24VT.001, 111.24VT.002, and 112.25W1.001

@@ Abstract
We constrain the sky-projected obliquities of two low-density hot Neptune planets, HATS-38 b and WASP-139 b, orbiting nearby G and K stars using Rossiter-McLaughlin (RM) observations with VLT/ESPRESSO, yielding $\lambda=-108_{-16}^{+11}$ deg and $-85.6_{-4.2}^{+7.7}$ deg, respectively. To model the RM effect, we use a new publicly available code, ironman, which is capable of jointly fitting transit photometry, Keplerian radial velocities, and RM effects. WASP-139 b has a residual eccentricity $e=0.103_{-0.041}^{+0.050}$ while HATS-38 b has an eccentricity of $e=0.112_{-0.070}^{+0.072}$, which is compatible with a circular orbit given our data. Using the obliquity constraints, we show that they join a growing group of hot and low-density Neptunes on polar orbits. We use long-term radial velocities to rule out companions with masses $\sim 0.3-50$ $M_{J}$ within $\sim 10$ au. We show that the orbital architectures of the two Neptunes can be explained with high-eccentricity migration from $\gtrsim 2$ au driven by an unseen distant companion. If HATS-38b has no residual eccentricity, its polar and circular orbit can also be consistent with a primordial misalignment. Finally, we performed a hierarchical Bayesian modeling of the true obliquity distribution of Neptunes and found suggestive evidence for a higher preponderance of polar orbits of hot Neptunes compared to Jupiters. However, we note that the exact distribution is sensitive to the choice of priors, highlighting the need for additional obliquity measurements of Neptunes to robustly compare the hot Neptune obliquity distribution to Jupiters.
@@ 1 Introduction
However, the signature of the dynamical evolution is the true 3D obliquity $\psi$, not its sky projection $\lambda$. The true obliquity $\psi$ can be estimated from $\lambda$ measured by the RM effect through combining it with constraints on the orbital inclination, $i$ constrained by the shape of a planetary transit, and the stellar inclination, $i_{\star}$, which can be constrained by measuring the stellar rotation period ($P_{\rm rot}$), the projected rotational velocity ($v\sin i_{\star}$), and the radius of the host star (e.g., Hirano et al., 2014; Morton & Winn, 2014; Masuda & Winn, 2020).
@@
Thanks to recent observational efforts, the current sample of planets with measured obliquity shows a growing number of smaller planets including hot Neptunes ($5010<M_{p}/M_{\oplus}<$ 50 or $62<R_{p}/R_{\oplus}<$ 6) in the “hot Neptune desert”, a region of the period-radius diagram where Neptune-sized exoplanets are notably rare (Mazeh et al., 2016) likely due to intense stellar radiation causing atmospheric erosion. Interestingly, many hot Neptunes are observed to be in close-to-polar orbits and have evaporating atmospheres, including HAT-P-11 b (Sanchis-Ojeda & Winn, 2011), GJ 436 b (Bourrier et al., 2018, 2022), WASP-107 b (Dai & Winn, 2017; Rubenzahl et al., 2021), and GJ 3470 b (Stefansson et al., 2022), suggesting a possible link between the Neptune desert and the dynamical history of low-mass gas giants. Theoretical scenarios that are capable of explaining the polar orbits of these planets require the presence of massive distant companions (e.g., Petrovich et al., 2020). Recently, outer companions have been confirmed in the HAT-P-11 (Yee et al., 2018) and WASP-107 (Piaulet et al., 2021) systems, suggesting that the transiting planets arrived at their current polar orbits through dynamical interactions (see Lu et al., 2024; Yu & Dai, 2024). As for the atmospheres, the orbital eccentricities may be large enough so that tidal heating can inflate these planets (e.g., by a factor of $\simeq 1.5-1.7$ for GJ 3470 b; Stefansson et al., 2022), possibly accounting for their evaporation. Nevertheless, the sample of hot Neptunes with measured obliquities is still small, and more measurements are necessary to test these emerging trends and asses the underlying mechanism behind the formation of this transiting population.
In this work, we present ESPRESSO (Pepe et al., 2021) observations of the RM effect of the two hot Neptunes HATS-38 b (Jordán et al., 2020) and WASP-139 b (Hellier et al., 2017). HATS-38 b is a hot Neptune in the middle of the hot Neptune desert orbiting a G-dwarf host star, while WASP-139 b is a super-Neptune orbiting a K-star. The precise in-transit spectroscopic observations with ESPRESSO reveal that both planets are in nearly polar orbits, continuing with the observed trends for low-density gas giants. We also performed a long-term radial velocity (RV) monitoring of both stars in order to constrain the presence of distant planetary companions in the systems.
@@
CHAT 2017 Feb 05 $r^{\prime}$ filter Jordán et al. (2020)
@@ 2.3 Observatoire Moana Photometry
Simultaneously with ESPRESSO observations, we observed the transit of HATS-38 b using the station of the Observatoire Moana located in El Sauce (ES) Observatory in Chile. Observatoire Moana is a global network of small-aperture robotic optical telescopes. The ES station consists of a 0.6 m CDK robotic telescope coupled to an Andor iKon-L 936 deep depletion 2k $\times$ 2k CCD with a scale of 0.67′′ per pixel. For these observations, we used a Sloan $r^{\prime}$ filter, and the exposure time was set to 74 s. Data reduction was done using a dedicated pipeline that automatically performs the CCD reduction steps, followed by the measurement of the aperture photometry for the brightest stars in the field. The pipeline also generates the differential light curve of the target star by identifying the optimal comparison stars based on color, brightness, and proximity to the target. The MOANA/ES light curve is shown in panel h of Figure 1, along with the best model.
@@ 3 Stellar Parameters
The second step consists of obtaining the physical parameters of the star by using publicly available broad-band magnitudes of the star and comparing them with those produced by different PARSEC stellar evolutionary models (Bressan et al., 2012) by taking into account the distance to the star computed from the Gaia DR2 (Gaia Collaboration et al., 2018) parallax. This procedure delivers a new value of $\log{g}$ that is held fixed in a new run of zaspe. We iterate between the two procedures until reaching convergence, which happens when two consecutive zaspe runs deliver the same values of $T_{\rm eff}$ and [Fe/H]222these parameters come from a grid search, so that for small changes in $\log{g}$ ($\Delta\log{g}<0.01$) the model with the best matching $T_{\rm eff}$ and [Fe/H] will be the same. The obtained parameters for each star are presented in Table 2.
@@ 5 Joint fit
Figure 3: a) Sky-projected obliquities of Neptunes ($5010<M_{p}/M_{\oplus}<$ 50 or $62<R_{p}/R_{\oplus}<$ 6) as a function of stellar effective temperature (blue). Neptunes on compact multiplanet systems are shown as squares, while isolated Neptunes are shown as circles. Jupiters (not Neptunes with $R_{p}/R_{\oplus}>6$) and super-Earths (not Neptunes with $R_{p}/R_{\oplus}<2$) are shown in orange and grey, respectively. Data from TEPCat (Southworth, 2011). b) Histograms of the $\lambda$ distribution for Neptunes (blue) and Jupiters (orange).
@@
With ironman, we analyzed both systems, including all the observations presented in Section 2. Given that we did not observe any significant variability in the TESS light curves that would allow us to measure the rotational period of the host stars, we have only fitted the sky-projected obliquity $\lambda$ and $v\sin{i_{\star}}$ here. For this analysis, we only considered the detrended TESS light curves (see Section 4) close to the transits, in windows of $\sim 6$ hours around the transit midpoint, to reduce the computational cost. To account for the instrumentation offsets and systematics, we included independent RV offsets and log-uniform jitter terms for each instrument. We placed uninformative priors for almost all parameters, except for the orbital periods and times of mid-transit, which had been constrained in Section 4, and for the stellar densities that were constrained in Section 3. Likewise for $\beta$ where we considered an instrumental broadening of 2.15 km/s because of the ESPRESSO resolution, and we used the macroturbulence law from Valenti & Fischer (2005), which yields a broadening of $\sim 4.4$ and $\sim 3.8$ km/s for $T_{\rm eff}=5662$ and $5233$ K, respectively. We added the instrumental and macroturbulence broadening in quadrature to set our priors for each target respectively, with an uncertainty of 2 km/s. The resulting posteriors of our fits for HATS-38 b and WASP-139 b are shown in Table 3.
Figures 1 and 2 show the different datasets together with the best-fit model. We found that both planets lie in a nearly-polar orbit, with sky-projected obliquities $\lambda=-108_{-16}^{+11}$ deg and $-85.6_{-4.2}^{+7.7}$ deg for HATS-38 b and WASP-139 b, respectively. Also, we found that both planets can be consistent with having eccentric orbits. This was not noted in Jordán et al. (2020) and Hellier et al. (2017), where they imposed circular orbits. In both cases, we tested fitting circular and eccentric orbits. For WASP-139 b the Bayesian evidence ($\log{Z}$) strongly favored the eccentric model, having a Bayes factor $\Delta\log{Z}=13.4$ and $\Delta{\rm AIC}=21.5$, so is the fit we report in Table 3. In contrast, for HATS-38 b, we have $\Delta\log{Z}=0.3$ and $\Delta{\rm AIC}=2.7$, suggesting that there is not sufficient evidence to favor the eccentric model. We note that the posteriors for all the parameters are consistent between the eccentric and circular fits, where the sky-projected obliquity is $\lambda=-111^{+12}_{-13}$ deg for the circular model in agreement with the value listed in Table 3. Although the circular fit formally has 2 fewer parameters than the eccentric fit, we elected to formally adopt the values from the eccentric fit to highlight the possible range of eccentricities compatible with the data. This possible eccentricity can be further constrained with additional RV or possibly secondary eclipse observations. Additionally, we also tested sampling the eccentricity and $\omega$ parameters as $\sqrt{e}\cos{\omega}$ and $\sqrt{e}\sin{\omega}$ which resulted in a set of parameters fully consistent with the ones reported in Table 3. We also tried fitting a long-term linear RV slope to the data, but it resulted in being consistent with zero so we fixed it to that value. Overall, the rest of the parameters are consistent with the ones reported in Jordán et al. (2020) and Hellier et al. (2017).
@@ 6 Discussion
In this work, we have characterized the orbits HATS-38 b and WASP-139 b, and found that they have nearly polar sky-projected obliquities, thus joining the growing sample of isolated (without nearby companions within 4 times the orbital period) low-density Neptunes in polar orbits that include HAT-P-11 b (Sanchis-Ojeda & Winn, 2011), WASP-107 b (Dai & Winn, 2017; Rubenzahl et al., 2021), GJ 436 b (Bourrier et al., 2018, 2022), and GJ 3470 b (Stefansson et al., 2022).
@@
$\beta_{\rm ESPRESSO}$ Intrinsic ESPRESSO stellar line width (km/s) $4.4_{-2.0}^{+1.9}$ $5.0_{-1.8}^{+1.8}$
@@
A related property of these planets is their residual eccentricities (potential eccentricity for HATS-38 b), which poses a problem for tidal circularization models, as these would predict circularization timescales generally shorter than their ages (see, e.g., Correia et al., 2020). Similarly to Espinoza-Retamal et al. (2023), we estimate the circularization timescale for HATS-38 b and WASP-139 b following Goldreich & Soter (1966). Assuming a modified quality factor (e.g., Ogilvie & Lin, 2007) $Q^{\prime}=10^{6}$ and pseudo-synchronous rotation of the planet, we get circularization timescales of $\sim 1.4$ and $\sim 2.3$ Gyr, for HATS-38 b and WASP-139 b respectively, which are timescales shorter than the estimated ages of the systems of $\sim 10$ and $\sim 6$ Gyr, continuing with the observed trend for polar Neptunes.
@@ 6.1 A possible preponderance of polar Neptunes?
To better quantify the possible preponderance of polar Neptunes in the sample, using the two new measurements we have provided here, we used a Hierarchical Bayesian model (HBM; e.g., Hogg et al., 2010) for the underlying obliquity distribution (e.g., Morton & Winn, 2014; Muñoz & Perets, 2018). In particular, we used the framework [^5] https://github.com/jiayindong/obliquity presented by Dong & Foreman-Mackey (2023), which models the underlying distribution of $\cos{\psi}$ across an exoplanet population using a mixture model of two Beta distributions (e.g., Gelman et al., 2014). This hierarchical Bayesian modeling framework allows the user to derive the true 3D obliquity distribution from observed sky-projected obliquities and has the capacity to capture anything from an isotropic distribution to a strongly bimodal population. In this work, we derived the $\cos{\psi}$ distribution only from the $\lambda$ measurements and did not include information about the stellar inclination.
@@
To derive the underlying $\cos{\psi}$ distribution, we ran two different hierarchical models with different priors on the compactness parameter $\kappa$: first, a model using informative priors on $\log\kappa$ following Dong & Foreman-Mackey (2023), and second, a fit using uninformative priors on $\log\kappa$. For the first model, we followed Dong & Foreman-Mackey (2023), and assumed an informative normal prior on $\log{\kappa}$ with a mean of 0 and a standard deviation of 3. Second, as an additional test—given the small size of the sample—we ran another model with an uninformative uniform prior between -4 and 10 for $\log{\kappa}$. In both cases, we used uniform priors on the location parameter $\mu$. The results of both of these model fits are shown in Table 4, and in Figure 4. For comparison, we also show in Figure 4 the distribution of the true 3D obliquity measurements from TEPCat (17 Neptunes and 30 Jupiters, Southworth, 2011).
@@
The observed difference in the obliquity distribution for the sample of Neptunes and Jupiters—with the former showing a possible preference for a polar + aligned obliquity dichotomy, while the latter shows a preference for an aligned + isotropic dichotomy—likely suggests that hot Jupiters and hot Neptunes could have formed through different formation channels. However, the sample of Neptunes is small compared to the one of Jupiters (27 versus 171 systems), and the exact parameters of the bimodal distribution are different with different priors. According to Dong & Foreman-Mackey (2023), in their tests, prior distribution choices did not strongly affect the derived distributions for samples of more than $\sim 50$ systems. Therefore, more measurements of the stellar obliquity of Neptunes will be required to further confirm or rule out these results.
Finally, we acknowledge that there are some biases that can affect these results given the small size of the Neptunes sample. Most of the Neptune obliquity measurements in our sample are measured via the RM effect (25 out of 27), except Kepler-30 b ($\lambda=4\pm 10$ deg, Sanchis-Ojeda et al., 2012) and TOI-3884 b ($\lambda=75\pm 10$ deg, Libby-Roberts et al., 2023) that were measured via starspost crossing events. As discussed in Siegel et al. (2023), the RM effect is intrinsically biased towards either well-aligned or retrograde orbits and is biased against the detection of polar orbits due to their lower RM amplitudes. This bias against polar orbits suggests that our result of detecting a candidate peak at polar orbits is robust—and, possibly, an underestimation. Future efforts like this one, and other ongoing programs like the Desert-Rim Exoplanets Atmosphere and Migration (DREAM, Bourrier et al., 2023), will be crucial for increasing the size of the sample and understanding the formation and evolution of the hot Neptunes.
Figure 4: Stellar obliquity distribution of Neptunes (blue) and Jupiters (orange). a) Inferred distribution using the HBM with an informative $\mathcal{N}(0,3)$ prior on the $\log{\kappa}$ parameter. This inference was done following the methodology from Dong & Foreman-Mackey (2023) and based purely on observed sky-projected obliquities available in the literature—combined with the two sky-projected obliquity measurements presented here—without including information about the stellar inclination. The lines and shaded regions represent the median and $1\sigma$ uncertainties of the inferred distributions. b) Same as panel a but using a uninformative $\mathcal{U}(-4,10)$ prior on $\log{\kappa}$. c) Observed true 3D stellar obliquities from TEPCat for visual reference (17 Neptunes and 30 Jupiters, Southworth, 2011). The true obliquity distribution also hints at a peak of polar Neptunes as inferred from the HBM modeling.